OpenStudy (anonymous):

Find all real number solutions for the equation. 5x 3 = 20x

OpenStudy (anonymous):

\[5x^3 = 20x\] \[5x^3 - 20x = 0\] 5x(x^2 - 4) = 0 Remember : \[a^2 - b^2 = (a+b)(a-b)\] Using this: \[5x(x+4)(x-4) = 0\] Either, x = 0 or (x + 4) = 0 or (x - 4) = 0 From here, can you solve it further??